OpenStudy (anonymous):

Can someone help me with 5 questions in geometry ??

OpenStudy (anonymous):

Shoot ahead.

OpenStudy (anonymous):

find the sum of the measures of the angles of the figure

OpenStudy (anonymous):

Assuming that this is a regular polygon-- the answer could be obtained by Sum of Interior Angles = (n-2) × 180°

OpenStudy (anonymous):

Where n is the number of angles.

OpenStudy (anonymous):

it's multiple choice : 900 1080 1620 1260

OpenStudy (anonymous):

Ok, so Answer= (7-2)*180.

OpenStudy (anonymous):

900?

OpenStudy (anonymous):

Yes. Good work.

OpenStudy (anonymous):

What is the sum of the angle measures of a 36-gon ?

OpenStudy (anonymous):

34*180=....

OpenStudy (anonymous):

http://www.mathsisfun.com/geometry/interior-angles-polygons.html This would help.

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

And next question....

OpenStudy (anonymous):

I finished the rest ! thanks for your help

OpenStudy (anonymous):

Cool. No prob.